# The completeness problem on 3-dimensional non-unimodular Lie groups

### v2

@@ 1. Introduction
Any simply connected solvable Lie group $G$ is given by a semi-direct product $\mathbb{R}\ltimes_{A}\mathbb{R}^{2}$, where $A$ is a $2\times 2$ real matrix and $\mathbb{R}$ acts on $\mathbb{R}^{2}$ via $t\mapsto\mathrm{exp}\,tA$. The isomorphism class of the Lie algebra $\mathfrak{g}$ of $G$ is thus determined by the Jordan normal form of $A$. We thus have the following types of solvable Lie algebras, cf. [Bia98], (i) $A$ is not diagonalizable over $\mathbb{C}$, (ii) $A$ is diagonalizable with real eigenvalues or (iii) $A$ is diagonalizable over $\mathbb{C}$ with complex non-real eigenvalues.
There is, up to isomorphism, only one non-unimodular Lie algebra of type (i), the Lie algebra which we denote by $\mathfrak{psh}$. This case was considered in the first part of our study [CFZ24]. More precisely, we solved the completeness problem for the pseudo-homothetic Lie group, the simply connected Lie group whose Lie algebra is $\mathfrak{psh}$. In this manuscript, we investigate the case where $A$ is diagonalizable with real eigenvalues. Up to normalization with respect to the eigenvalue of highest absolute value, $A$ is similar to $\mathrm{diag}(1,\lambda)$, with $|\lambda|\leq 1$. We thus obtain a 1-parameter family of isomorphism classes of Lie algebras $\mathfrak{h}(\lambda)$, $|\lambda|\leq 1$, such that the non-trivial bracket relations on $\mathfrak{h}(\lambda)$ are given by
$[e_{1},e_{2}]=e_{2},\quad[e_{1},e_{3}]=\lambda e_{3}.$ (1.1)
@@
The only unimodular Lie algebra in this family is obtained for the limiting case $\lambda=-1$ and corresponds to the Lie algebra of the Poincaré group (also called $\mathrm{Sol}$), the group of isometries of the 2-dimensional Minkowski space. Another limiting case occurs when $\lambda=1$ for the 3-dimensional homothetic group $\mathrm{Ho}$ given by $\mathbb{R}\ltimes_{\mathrm{Id}}\mathbb{R}^{2}$. It was proven in [Gue96], that all left-invariant Lorentzian metrics on homothetic groups $\mathbb{R}\ltimes_{\mathrm{Id}}\mathbb{R}^{n}$ of any dimension are incomplete. Moreover, the rather remarkable fact that all indefinite left-invariant metrics on $\mathbb{R}\ltimes_{\mathrm{Id}}\mathbb{R}^{n}$ are incomplete was later extended in [VS20].
@@
We briefly recall that the study of the geodesic flow of a left-invariant metric on a Lie group $G$ can be reduced to the study of the flow of a certain vector field on $\mathfrak{g}$, which we shall call the geodesic field, via the Euler-Arnold formalism [Arn66]. We refer to [CFZ24, Section 2] for an account of background material and techniques on this topic.
With this formalism in mind and for both easiness and clearness of exposition, we will, throughout the text, be referring to geodesic completeness of metric Lie algebras $(\mathfrak{g},q)$ to mean the geodesic completeness of $(G,g)$ where $G$ is a Lie group with Lie algebra $\mathfrak{g}$ and $g$ is the left-invariant metric defined by $q$.
@@
We proceed, in Sec. 5, with the study of geodesic completeness on the Lie algebras $\mathfrak{h}(\lambda)$ for $10<|\lambda|<$ 1. Taking $B=\{e_{1},e_{2},e_{3}\}$ to be a standard basis of $\mathfrak{h}(\lambda)$, we establish the following characterization.
@@
Let $\mathfrak{h}(\lambda)$, $10<|\lambda|<$ 1, be equipped with a Lorentzian metric $q$. Then, $q$ is geodesically complete if and only if $e_{3}$ is timelike and $e_{2}$ is not spacelike. Furthermore, if $e_{3}$ is timelike and
@@
In Sec. 6, we quickly study the limiting Lie algebras $\mathfrak{h}(\lambda)$, $\lambda=\pm 1$. We recover the result of [Gue96] for $\mathfrak{h}(1)$ and clarify a misstatement in [BM08, Prop. 3] for $\mathfrak{h}(-1)$.
We finish this paper, with Sec. 7, by briefly considering the remaining family of non-unimodular Lie algebras of case (iii), in order to prove the following relevant result.
@@
The classification of normal forms of left-invariant Lorentzian metrics in dimension 3 has been considered more or less implicitly in several articles. A detailed account was presented in [HL23], see also Appendix A in [CFZ25]. We include some computations here, for 3-dimensional Lie algebras $\mathfrak{h}(\lambda)$ where $|\lambda|<1$. Consider a standard basis $B=\{e_{1},e_{2},e_{3}\}$ satisfying the bracket relations
$[e_{1},e_{2}]=e_{2},\ [e_{1},e_{3}]=\lambda e_{3},\ [e_{2},e_{3}]=0.$
@@
As can be seen, $\mathrm{Aut}(\mathfrak{h}(\lambda))$ is 4-dimensional and has four connected components with $\mathrm{Aut}(\mathfrak{h}(\lambda))/\mathrm{Aut}_{0}(\mathfrak{h}(\lambda))$ isomorphic to the Klein group $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$. Consider the generic $3\times 3$ matrix
@@
which is assumed to represent a non-degenerate symmetric bilinear form on the basis $B$. The image of $m$ under the automorphism
@@
$\begin{array}[]{llll}\mathscr{Q}_{0}=\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix},&\mathscr{Q}_{1}=\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&-1\end{pmatrix},&\mathscr{Q}_{2}=\begin{pmatrix}1&0&0\\ 0&-1&0\\ 0&0&1\end{pmatrix},&\mathscr{Q}_{3}=\begin{pmatrix}-1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix}.\end{array}$
@@
$\begin{array}[]{lll}\mathscr{Q}_{4,r}=\begin{pmatrix}1&0&0\\ 0&1&1\\ 0&1&r\end{pmatrix},&\mathscr{Q}_{5,s}=\begin{pmatrix}1&0&0\\ 0&-1&1\\ 0&1&s\end{pmatrix},&\mathscr{Q}_{6,t}=\begin{pmatrix}-1&0&0\\ 0&1&1\\ 0&1&t\end{pmatrix},\end{array}$
@@
Although not our main object of study, we include here the metric normal form for $\mathfrak{h}(\lambda),\lambda=\pm 1$, since they can be easily computed and will be useful in Sec. 6.
The automophism group of $\mathfrak{h}(-1)$ is given by
@@
with $r\neq 1$, $-1<s\leq 0$ and $t>1$.
The automophism group of $\mathfrak{h}(1)$ is given by
@@
 bilinear form  geodesic field  invariant planes  idempotents $\mathscr{Q}_{1}$ $\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&-1\end{pmatrix}$$\begin{cases}\dot{x}=-y^{2}+\lambda z^{2}\\ \dot{y}=xy\\ \dot{z}=\lambda xz\end{cases}$ $y=0$ $z=0$ Yes, if $\lambda\neq 0$ $\mathscr{Q}_{2}$ $\begin{pmatrix}1&0&0\\ 0&-1&0\\ 0&0&1\end{pmatrix}$$\begin{cases}\dot{x}=y^{2}-\lambda z^{2}\\ \dot{y}=xy\\ \dot{z}=\lambda xz\end{cases}$ $y=0$ $z=0$ Yes $\mathscr{Q}_{3}$ $\begin{pmatrix}-1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix}$$\begin{cases}\dot{x}=y^{2}+\lambda z^{2}\\ \dot{y}=xy\\ \dot{z}=\lambda xz\end{cases}$ $y=0$ $z=0$ Yes $\mathscr{Q}_{4,r}$ $r<1$ $\begin{pmatrix}1&0&0\\ 0&1&1\\ 0&1&r\end{pmatrix}$$\begin{cases}\dot{x}=-y^{2}-\lambda rz^{2}-(\lambda+1)yz\\ \dot{y}=\frac{x}{r-1}((r-\lambda)y+r(1-\lambda)z)\\ \dot{z}=\frac{x}{r-1}((\lambda-1)y+(r\lambda-1)z)\end{cases}$ $y=-z$ $y=-rz$ Yes, if $10<r<$ 1 or $\lambda\neq 0$ and $r<1$ $\mathscr{Q}_{5,s}$ $s>-1$ $\begin{pmatrix}1&0&0\\ 0&-1&1\\ 0&1&s\end{pmatrix}$$\begin{cases}\dot{x}=y^{2}-\lambda sz^{2}-(\lambda+1)yz\\ \dot{y}=\frac{x}{s+1}((s+\lambda)y+s(\lambda-1)z)\\ \dot{z}=\frac{x}{s+1}((\lambda-1)y+(s\lambda+1)z)\end{cases}$ $y=z$ $y=-sz$ Yes, if $s>0$ $\mathscr{Q}_{6,t}$ $t>1$ $\begin{pmatrix}-1&0&0\\ 0&1&1\\ 0&1&t\end{pmatrix}$$\begin{cases}\dot{x}=y^{2}+\lambda tz^{2}+(\lambda+1)yz\\ \dot{y}=\frac{x}{t-1}((t-\lambda)y+t(1-\lambda)z)\\ \dot{z}=\frac{x}{t-1}((\lambda-1)y+(t\lambda-1)z)\end{cases}$ $y=-z$ $y=-tz$ Yes $\mathscr{Q}_{7}$ $\begin{pmatrix}1&0&0\\ 0&0&1\\ 0&1&1\end{pmatrix}$$\begin{cases}\dot{x}=-\lambda z^{2}-(\lambda+1)yz\\ \dot{y}=\lambda xy+(\lambda-1)xz\\ \dot{z}=xz\end{cases}$ $y=-z$ $z=0$ Yes $\mathscr{Q}_{8}$ $\begin{pmatrix}1&0&0\\ 0&0&1\\ 0&1&-1\end{pmatrix}$$\begin{cases}\dot{x}=\lambda z^{2}-(\lambda+1)yz\\ \dot{y}=\lambda xy+(1-\lambda)xz\\ \dot{z}=xz\end{cases}$ $y=z$ $z=0$ No $\mathscr{Q}_{9}$ $\begin{pmatrix}1&0&0\\ 0&0&1\\ 0&1&0\end{pmatrix}$$\begin{cases}\dot{x}=-(\lambda+1)yz\\ \dot{y}=\lambda xy\\ \dot{z}=xz\end{cases}$ $y=0$ $z=0$ No $\mathscr{Q}_{10}$ $\begin{pmatrix}0&0&1\\ 0&1&1\\ 1&1&1\end{pmatrix}$$\begin{cases}\dot{x}=\lambda x^{2}+(\lambda-1)(y+z)x\\ \dot{y}=y^{2}+\lambda z^{2}+xy+(\lambda+1)(x+y)z\\ \dot{z}=-y^{2}-\lambda z^{2}-(\lambda x+(1+\lambda)y)z\end{cases}$ $x=0$ $y=-z$ Yes $\mathscr{Q}_{11}$ $\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&1\end{pmatrix}$$\begin{cases}\dot{x}=x^{2}\\ \dot{y}=-\lambda z^{2}-xy\\ \dot{z}=\lambda xz\end{cases}$ $x=0$ $z=0$ Yes $\mathscr{Q}_{12}$ $\begin{pmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{pmatrix}$$\begin{cases}\dot{x}=\lambda x^{2}\\ \dot{y}=xy\\ \dot{z}=-y^{2}-\lambda xz\end{cases}$ $x=0$ $y=0$ Yes, if $\lambda\neq 0$
@@
In Table 1, we listed two invariant planes for each geodesic vector field. We note that, for $10<|\lambda|<$ 1, no further invariant planes exist. However, this is not the case for $\lambda=0$, where other invariant planes do exist. For instance, for $\mathscr{Q}_{3}$ infinitely many invariant planes exist, namely $z=c$, for every $c\in\mathbb{R}$. Another example is that of $\mathscr{Q}_{7}$ where $y=-x$ is an additional invariant plane. When the system of differential equations does not look tractable, the existence of invariant planes can be an important tool in the study of the vector field dynamics, by means of a linear change of variables. See Subsec. 5.2.2 for an example. Notice, however, that a geodesic field may not have invariant planes, as is the case of (7.1).
@@
Let $\mathfrak{z}$ be the center and $\mathfrak{d}$ be the derived subalgebra of $\mathfrak{aff}(\mathbb{R})\oplus\mathbb{R}$. In our standard basis, the non-zero bracket relation is given by $[e_{1},e_{2}]=e_{2}$, and thus $\mathfrak{z}=\mathrm{span}\{e_{3}\}$ and $\mathfrak{d}=\mathrm{span}\{e_{2}\}$. We start by observing that both $\mathfrak{z}$ and $\mathfrak{d}$ are invariant under the action of $\mathrm{Aut}(\mathfrak{aff}(\mathbb{R})\oplus\mathbb{R})$ and that the induced action on the space of Lorentzian metrics preserves their type (spacelike, timelike or lightlike) as well as whether they are orthogonal or not.
@@
  • $\mathfrak{z}$ is spacelike: $\mathscr{Q}_{2}$, $\mathscr{Q}_{3}$, $\mathscr{Q}_{4,r}$ ($10<r<$ 1), $\mathscr{Q}_{5,s}$ ($s>0$), $\mathscr{Q}_{6,t}$, $\mathscr{Q}_{7}$, $\mathscr{Q}_{10}$, $\mathscr{Q}_{11}$;
  • $\mathfrak{z}$ is timelike: $\mathscr{Q}_{1}$, $\mathscr{Q}_{4,r}$ ($r<0$), $\mathscr{Q}_{5,s}$ ($-1<s<0$), $\mathscr{Q}_{8}$;

### v3

@@ 1. Introduction
Any simply connected solvable Lie group $G$ is given by a semidirect product $\mathbb{R}\ltimes_{A}\mathbb{R}^{2}$, where $A$ is a $2\times 2$ real matrix and $\mathbb{R}$ acts on $\mathbb{R}^{2}$ via $t\mapsto\mathrm{exp}\,tA$. The isomorphism class of the Lie algebra $\mathfrak{g}$ of $G$ is thus determined by the Jordan normal form of $A$. We therefore have the following types of solvable Lie algebras, cf. [Bia98], (i) $A$ is not diagonalizable over $\mathbb{C}$, (ii) $A$ is diagonalizable with real eigenvalues or (iii) $A$ is diagonalizable over $\mathbb{C}$ with complex non-real eigenvalues.
There is, up to isomorphism, only one non-unimodular Lie algebra of type (i), the Lie algebra which we denote by $\mathfrak{psh}$. This case was considered in the first part of our study [CFZ25a]. More precisely, we solved the completeness problem for the pseudo-homothetic Lie group, the simply connected Lie group whose Lie algebra is $\mathfrak{psh}$. In this manuscript, we investigate the case where $A$ is diagonalizable with real eigenvalues. Up to normalization with respect to the eigenvalue of highest absolute value, $A$ is similar to $\mathrm{diag}(1,\lambda)$, with $|\lambda|\leq 1$. We thus obtain a 1-parameter family of isomorphism classes of Lie algebras $\mathfrak{h}(\lambda)$, $|\lambda|\leq 1$, such that the non-trivial bracket relations on $\mathfrak{h}(\lambda)$ are given by
$e_{1},e_{2}]=e_{2},\quad[e_{1},e_{3}]=\lambda e_{3}.$ (1.1)
@@
The only unimodular Lie algebra in this family is obtained for the limiting case $\lambda=-1$ and corresponds to the Lie algebra of the Poincaré group (also called $\mathrm{Sol}$), the group of isometries of the 2-dimensional Minkowski space. Another limiting case occurs when $\lambda=1$ for the 3-dimensional homothetic group $\mathrm{Ho}$ given by $\mathbb{R}\ltimes_{\mathrm{Id}}\mathbb{R}^{2}$. It was proven in [Gue96] that all left-invariant Lorentzian metrics on homothetic groups $\mathbb{R}\ltimes_{\mathrm{Id}}\mathbb{R}^{n}$ of any dimension are incomplete. Moreover, the rather remarkable fact that all indefinite left-invariant metrics on $\mathbb{R}\ltimes_{\mathrm{Id}}\mathbb{R}^{n}$ are incomplete was later extended in [VS20].
@@
A classical and fundamental fact is that the study of the geodesic flow of a left-invariant metric on a Lie group $G$ can be reduced to the study of the flow of a certain vector field on $\mathfrak{g}$, which we shall call the geodesic field, via the Euler-Arnold formalism [Arn66]. We present in Sec. 2 a brief account of background material and techniques on this topic. With this formalism in mind and for ease of exposition, we will speak of the geodesic completeness of a metric Lie algebra $(\mathfrak{g},q)$ to mean the geodesic completeness of $(G,g)$, where $G$ is a Lie group with Lie algebra $\mathfrak{g}$ and $g$ is the left-invariant metric defined by $q$.
@@
We proceed, in Sec. 6, with the study of geodesic completeness on the Lie algebras $\mathfrak{h}(\lambda)$ for $0<|\lambda|<1$. Taking $B=\{e_{1},e_{2},e_{3}\}$ to be a standard basis of $\mathfrak{h}(\lambda)$, we establish the following characterization.
@@
Let $\mathfrak{h}(\lambda)$, $0<|\lambda|<1$, be equipped with a Lorentzian metric $q$. Then, $q$ is geodesically complete if and only if $e_{3}$ is timelike and $e_{2}$ is not spacelike. Furthermore, if $e_{3}$ is timelike and
@@
In Sec. 7, we study the limiting Lie algebras $\mathfrak{h}(\lambda)$, $\lambda=\pm 1$. We recover the result of [Gue96] for $\mathfrak{h}(1)$ and clarify a misstatement in [BM08, Prop. 3] for $\mathfrak{h}(-1)$.
We finish this paper with Sec. 8, by briefly considering the remaining family of non-unimodular Lie algebras of case (iii), in order to prove the following relevant result.
@@
2. Preliminaries
In this section, we briefly recall the basic framework used throughout the text, fixing notation and terminology. An extended exposition can be found in [CFZ25a, Sec. 2] and references therein.
Let us fix a Lie group $G$ with Lie algebra $\mathfrak{g}=T_{e}G$. Left-invariant metrics on $G$ are in one-to-one correspondence with non-degenerate symmetric bilinear forms on $\mathfrak{g}$. The Euler-Arnold theorem [Arn66] establishes that the geodesic flow of $(G,g)$ is uniquely associated with the flow of a vector field on $(\mathfrak{g},q=g_{e})$ as follows. If $\gamma$ is a geodesic with $\gamma(0)=e$, its velocity $\dot{\gamma}$ can be identified with a curve $v$ in $\mathfrak{g}$ via left translation. The geodesic equation then reduces to
$\dot{v}(t)=\operatorname{ad}^{\dagger}_{v(t)}v(t),$ (2.1)
where $\operatorname{ad}^{\dagger}$ denotes the $q$-transpose of the adjoint map $\operatorname{ad}$. For simplicity of language, we will refer to the vector field defined by the system of ODEs 2.1 as the geodesic field. Clearly, a left-invariant metric on $G$ is geodesically complete if and only if its associated geodesic field on $\mathfrak{g}$ is complete.
The geodesic field admits a natural first integral: the energy associated with $q$ is conserved along trajectories. Further first integrals (eventually non-polynomial) may occur, providing effective constraints on the flow. For instance, if a positive-definite quadratic first integral exists then the flow is complete. On the other hand, a very useful tool for proving incompleteness is given by idempotents: an element $v_{o}\in\mathfrak{g}$ is called an idempotent if the geodesic field at $v_{o}$ is equal to $v_{o}$. The existence of an idempotent leads to an integral curve that blows up in finite time. However, not all incomplete trajectories arise from idempotents. The following lemma, [CFZ25a, Lemma 2.3], will be instrumental in providing such curves.
Lemma 2.1.
Let $(E)$ be an ordinary differential equation of the form
$\dot{x}(t)=ax^{2}(t)+\alpha(t),$
such that $a>0$ and $\alpha\in\mathscr{C}^{\infty}(\mathbb{R})$; $t\mapsto\alpha(t)\geq 0$. Let $\gamma\colon I\rightarrow\mathbb{R}$ be a nonzero maximal integral curve of $(E)$, then $\gamma$ must be incomplete.
Finally, another powerful tool is the automorphism group of the Lie algebra $\mathfrak{g}$, $\operatorname{Aut}(\mathfrak{g})$. The automorphism group acts naturally on $\mathrm{Sym}^{\ast}(\mathfrak{g})$, the space of non-degenerate symmetric bilinear forms on $\mathfrak{g}$. Not surprisingly, completeness of the flow of the geodesic field of $(\mathfrak{g},q)$ is invariant under this action. This was proved, for instance, in [EFSZ24]. Moreover, the geodesic flow remains unchanged under rescaling. Therefore, all bilinear forms in each orbit of $\mathrm{Sym}^{\ast}(\mathfrak{g})$ by the action of $\mathbb{R}^{\ast}\times\mathrm{Aut}(\mathfrak{g})$ are either complete or incomplete. Representatives of such orbits are usually called metric normal forms.
@@ 2. Preliminaries
The classification of normal forms of left-invariant Lorentzian metrics in dimension 3 has been considered more or less implicitly in several articles. A detailed account was presented in [HL23], see also Appendix A in [CFZ25b]. We include some computations here, for 3-dimensional Lie algebras $\mathfrak{h}(\lambda)$ where $|\lambda|<1$. Consider a standard basis $B=\{e_{1},e_{2},e_{3}\}$ satisfying the bracket relations
$e_{1},e_{2}]=e_{2},\ [e_{1},e_{3}]=\lambda e_{3},\ [e_{2},e_{3}]=0.$
@@
As can be seen, $\mathrm{Aut}(\mathfrak{h}(\lambda))$ is 4-dimensional and has four connected components, with $\mathrm{Aut}(\mathfrak{h}(\lambda))/\mathrm{Aut}^{0}(\mathfrak{h}(\lambda))$ isomorphic to the Klein group $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$. Consider the generic $3\times 3$ matrix
@@
which is assumed to represent a non-degenerate symmetric bilinear form with respect the basis $B$. The image of $m$ under the automorphism
@@
[…]
@@
[…]
@@
Although not our main object of study, we include here the metric normal forms for $\mathfrak{h}(\lambda),\lambda=\pm 1$, since they can be easily computed and will be useful in Sec. 7.
The automorphism group of $\mathfrak{h}(-1)$ is given by
@@
with $r\neq 1$, $1<s\leq 0$ and $t>1$.
The automorphism group of $\mathfrak{h}(1)$ is given by
@@
 bilinear form  geodesic field  invariant planes  idempotents $\mathscr{Q}_{1}$ $\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&-1\end{pmatrix}$ $\begin{cases}\dot{x}=-y^{2}+\lambda z^{2}\\ \dot{y}=xy\\ \dot{z}=\lambda xz\end{cases}$ $y=0$ $z=0$ Yes, if $\lambda\neq 0$ $\mathscr{Q}_{2}$ $\begin{pmatrix}1&0&0\\ 0&-1&0\\ 0&0&1\end{pmatrix}$ $\begin{cases}\dot{x}=y^{2}-\lambda z^{2}\\ \dot{y}=xy\\ \dot{z}=\lambda xz\end{cases}$ $y=0$ $z=0$ Yes $\mathscr{Q}_{3}$ $\begin{pmatrix}-1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix}$ $\begin{cases}\dot{x}=y^{2}+\lambda z^{2}\\ \dot{y}=xy\\ \dot{z}=\lambda xz\end{cases}$ $y=0$ $z=0$ Yes $\mathscr{Q}_{4,r}$ $r<1$ $\begin{pmatrix}1&0&0\\ 0&1&1\\ 0&1&r\end{pmatrix}$ $\begin{cases}\dot{x}=-y^{2}-\lambda rz^{2}-(\lambda+1)yz\\ \dot{y}=\frac{x}{r-1}((r-\lambda)y+r(1-\lambda)z)\\ \dot{z}=\frac{x}{r-1}((\lambda-1)y+(r\lambda-1)z)\end{cases}$ y=zy=-z y=rzy=-rz Yes, if $0<r<1$ or $\lambda\neq 0$ and $r<1$ $\mathscr{Q}_{5,s}$ s>1s>-1 $\begin{pmatrix}1&0&0\\ 0&-1&1\\ 0&1&s\end{pmatrix}$ $\begin{cases}\dot{x}=y^{2}-\lambda sz^{2}-(\lambda+1)yz\\ \dot{y}=\frac{x}{s+1}((s+\lambda)y+s(\lambda-1)z)\\ \dot{z}=\frac{x}{s+1}((\lambda-1)y+(s\lambda+1)z)\end{cases}$ $y=z$ y=szy=-sz Yes, if $s>0$ $\mathscr{Q}_{6,t}$ $t>1$ $\begin{pmatrix}-1&0&0\\ 0&1&1\\ 0&1&t\end{pmatrix}$ $\begin{cases}\dot{x}=y^{2}+\lambda tz^{2}+(\lambda+1)yz\\ \dot{y}=\frac{x}{t-1}((t-\lambda)y+t(1-\lambda)z)\\ \dot{z}=\frac{x}{t-1}((\lambda-1)y+(t\lambda-1)z)\end{cases}$ y=zy=-z y=tzy=-tz Yes $\mathscr{Q}_{7}$ $\begin{pmatrix}1&0&0\\ 0&0&1\\ 0&1&1\end{pmatrix}$ $\begin{cases}\dot{x}=-\lambda z^{2}-(\lambda+1)yz\\ \dot{y}=\lambda xy+(\lambda-1)xz\\ \dot{z}=xz\end{cases}$ y=zy=-z $z=0$ Yes $\mathscr{Q}_{8}$ $\begin{pmatrix}1&0&0\\ 0&0&1\\ 0&1&-1\end{pmatrix}$ $\begin{cases}\dot{x}=\lambda z^{2}-(\lambda+1)yz\\ \dot{y}=\lambda xy+(1-\lambda)xz\\ \dot{z}=xz\end{cases}$ $y=z$ $z=0$ No $\mathscr{Q}_{9}$ $\begin{pmatrix}1&0&0\\ 0&0&1\\ 0&1&0\end{pmatrix}$ $\begin{cases}\dot{x}=-(\lambda+1)yz\\ \dot{y}=\lambda xy\\ \dot{z}=xz\end{cases}$ $y=0$ $z=0$ No $\mathscr{Q}_{10}$ $\begin{pmatrix}0&0&1\\ 0&1&1\\ 1&1&1\end{pmatrix}$ $\begin{cases}\dot{x}=\lambda x^{2}+(\lambda-1)(y+z)x\\ \dot{y}=y^{2}+\lambda z^{2}+xy+(\lambda+1)(x+y)z\\ \dot{z}=-y^{2}-\lambda z^{2}-(\lambda x+(1+\lambda)y)z\end{cases}$ $x=0$ y=zy=-z Yes $\mathscr{Q}_{11}$ $\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&1\end{pmatrix}$ $\begin{cases}\dot{x}=x^{2}\\ \dot{y}=-\lambda z^{2}-xy\\ \dot{z}=\lambda xz\end{cases}$ $x=0$ $z=0$ Yes $\mathscr{Q}_{12}$ $\begin{pmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{pmatrix}$ $\begin{cases}\dot{x}=\lambda x^{2}\\ \dot{y}=xy\\ \dot{z}=-y^{2}-\lambda xz\end{cases}$ $x=0$ $y=0$ Yes, if $\lambda\neq 0$
@@
In Table 1, we listed two invariant planes for each geodesic vector field. We note that, for $0<|\lambda|<1$, no further invariant planes exist. However, this is not the case for $\lambda=0$, where other invariant planes do exist. For instance, for $\mathscr{Q}_{3}$, infinitely many invariant planes exist, namely $z=c$, for every $c\in\mathbb{R}$. Another example is that of $\mathscr{Q}_{7}$, where y=xy=-x is an additional invariant plane. When the system of differential equations does not look tractable, the existence of invariant planes can be an important tool in the study of the vector field dynamics, by means of a linear change of variables. See Subsec. 6.2.2 for an example. Notice, however, that a geodesic field may not have invariant planes, as is the case of (8.1).
@@
Let $\mathfrak{z}$ be the center and $\mathfrak{d}$ be the derived subalgebra of $\mathfrak{aff}(\mathbb{R})\oplus\mathbb{R}$. In our standard basis, the non-zero bracket relation is given by $e_{1},e_{2}]=e_{2}$, and thus $\mathfrak{z}=\mathrm{span}\{e_{3}\}$ and $\mathfrak{d}=\mathrm{span}\{e_{2}\}$. We start by observing that both $\mathfrak{z}$ and $\mathfrak{d}$ are invariant under the action of $\mathrm{Aut}(\mathfrak{aff}(\mathbb{R})\oplus\mathbb{R})$ and that the induced action on the space of Lorentzian metrics preserves their type (spacelike, timelike or lightlike) as well as whether they are orthogonal or not.
@@
  • $\mathfrak{z}$ is spacelike: $\mathscr{Q}_{2}$, $\mathscr{Q}_{3}$, $\mathscr{Q}_{4,r}$ ($0<r<1$), $\mathscr{Q}_{5,s}$ ($s>0$), $\mathscr{Q}_{6,t}$, $\mathscr{Q}_{7}$, $\mathscr{Q}_{10}$, $\mathscr{Q}_{11}$;
  • $\mathfrak{z}$ is timelike: $\mathscr{Q}_{1}$, $\mathscr{Q}_{4,r}$ ($r<0$), $\mathscr{Q}_{5,s}$ (1<s<0-1<s<0), $\mathscr{Q}_{8}$;